OpenStudy (anonymous):

Solve for y in the 1st equation. 7x +2y=19 x-y=4

OpenStudy (turingtest):

y=(1/2)(19-7x)

OpenStudy (turingtest):

is that all?

OpenStudy (anonymous):

I thought Y= 19-7x/ 2?

OpenStudy (anonymous):

where did you get the 1/2 from?

OpenStudy (turingtest):

7x +2y=19 2y=19-7x divide ALL by 2: y=(19-7x)/2=(1/2)(19-7x) if you just write Y= 19-7x/ 2 then you are not dividing the 19 by two, which you should be doing. Parentheses are important, keep track of them.

OpenStudy (anonymous):

sorry but I still don't understand how you got the 1/2

OpenStudy (turingtest):

do you understand that these two equations are the same?: y=(19-7x)/2 y=(1/2)(19-7x) In effect the WHOLE right side is being divided by 2 either way.

OpenStudy (turingtest):

here's another way to see it 2y=x multiply both sides by (1/2): y=(1/2)x=x/2 See? Dividing by 2 and multiplying by (1/2) is the same operation.

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

2y